# InverseMatrixVT3D: An Efficient Projection Matrix-Based Approach for 3D Occupancy Prediction

### v1

@@ Abstract
This paper introduces InverseMatrixVT3D, an efficient method for transforming multi-view image features into 3D feature volumes for 3D semantic occupancy prediction. Existing methods for constructing 3D volumes often rely on depth estimation, device-specific operators, or transformer queries, which hinders the widespread adoption of 3D occupancy models. In contrast, our approach leverages two projection matrices to store the static mapping relationships and matrix multiplications to efficiently generate global Bird’s Eye View (BEV) features and local 3D feature volumes. Specifically, we achieve this by performing matrix multiplications between multi-view image feature maps and two sparse projection matrices. We introduce a sparse matrix handling technique for the projection matrices to optimise GPU memory usage. Moreover, a global-local attention fusion module is proposed to integrate the global BEV features with the local 3D feature volumes to obtain the final 3D volume. We also employ a multi-scale supervision mechanism to further enhance performance. Comprehensive experiments on the nuScenes dataset demonstrate the simplicity and effectiveness of our method. The code will be made available at:https://github.com/DanielMing123/InverseMatrixVT3D
Figure 1: The figure above shows the pipeline of three approaches: Depth-Estimation-based approach (upper), Query-based approach (middle), and our approach (lower). We simplify the generation of the 3D Feature Volume by adopting a matrix multiplication method.
@@ I Introduction
Understanding the surrounding scene’s three-dimensional (3D) geometry is fundamental to developing autonomous driving (AV) systems. While lidar-based methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] which utilize explicit depth measurements have been performing exceptionally well on public datasets [14, 15, 16, 17], they are hindered by the expensive cost of sensors and the sparsity of data points. As a result, the broader application of lidar-based methods is limited.
@@
In recent years, vision-centric AV systems have garnered significant attention as a promising strategy due to its cost-effectiveness, stability, and generality. By utilizing multi-camera images as inputs, this approach has demonstrated competitive performance across various 3D perception tasks, including depth estimation [18, 19, 20], 3D object detection [21, 22, 23, 24, 25, 26, 27, 28], online high-definition (HD) map construction [29, 30, 31, 32, 33], and semantic map construction [34, 35, 36].
@@
3D object detection based on fusing surround view cameras can play a crucial role in 3D perception. However, it faces challenges in handling new scenarios. One of the challenges is the finite number of semantic classes in the training dataset, making it difficult to create a model for every potential scenario that might be encountered on the road. In contrast, a more practical approach to depict the vehicle’s surrounding environment is by directly reconstructing the 3D scenes. In pursuit of this objective, several methods [37, 38, 39, 40, 41] have been investigated to predict the 3D occupancy of a scene directly. These methods involve voxelizing the 3D space and assigning a probability to each voxel to determine its occupancy state–whether it is occupied or not. We argue that 3D occupancy serves as an adequate representation of the vehicle’s surrounding environment. This representation inherently ensures geometric consistency and can accurately describe occluded areas. Furthermore, it is more robust towards object classes that do not exist in the training dataset. Despite the promise of these methods, their inner structure is quite complex, and some methods require additional sensors to provide supervision signals. For instance, the method proposed in [40] relies on a lidar sensor to enhance performance through depth estimation supervision (Figure 1, upper). Additionally, the methods proposed in [38, 39] extensively employ query-based modules [42, 43, 44] to aggregate image features for the final 3D feature volume (Figure 1, middle).
@@
To efficiently and effectively represent a 3D scene using 3D occupancy, we propose InverseMatrixVT3D (Figure 1, bottom). Our method focuses on constructing projection matrices and simplifying the generation of local 3D feature volumes and global Bird’s Eye View (BEV) features through matrix multiplication between multi-scale feature maps and projection matrices. Additionally, we employ a sparse matrix handling technique to optimize GPU memory usage when using these sparse projection matrices. Furthermore, we introduce a global-local fusion module to integrate global BEV features with local 3D feature volumes, resulting in the final 3D volume. We also apply a multi-scale supervision mechanism to each level to further enhance performance. Through comparisons with state-of-the-art (SOTA) algorithms on the nuScenes benchmark, we demonstrate our method’s simplicity and superior overall performance.
@@
  • The novel projection matrix-based approach is proposed to construct local 3D feature volumes and global BEV features.
  • A global-local fusion module is proposed to integrate the information, resulting in the final 3D volume.
  • We compare our approach with other state-of-the-art (SOTA) algorithms in the 3D semantic occupancy prediction task to prove the superior performance of our method.
@@ II Related Work
II-A Depth-Estimation-Based BEV Perception
Lift-Splat-Shoot (LSS) [21] introduced a lift operation for depth estimation using multi-view cameras. This process generates a pseudo-3D point cloud by merging the pseudo-3D points obtained from the multi-view images. The pseudo-3D point cloud is projected onto the ego vehicle coordinate system by leveraging intrinsic and extrinsic parameters. Subsequently, a splat operation is employed to accumulate features in the same BEV plane grid with varying z-direction heights, resulting in the final BEV features. This study pioneered the generation of BEV features through depth estimation and the fusion of multi-camera information in BEV perception. Another important contribution in this area was made by BEVDepth [24], highlighting the crucial impact of depth estimation precision on the model’s performance. To address this, BEVDepth introduced a real lidar-sensor based 3D point cloud as an additional supervision signal for the depth estimation stage. Furthermore, BEVStereo [25] improved depth estimation performance by introducing the temporal stereo method.
@@ II-A Depth-Estimation-Based BEV Perception
II-B Depth-Estimation Based 3D Semantic Occupancy Prediction
Based on the success of depth-estimation based BEV perception algorithms, several works [45, 46, 41, 40] have focused on constructing the 3D feature volume using pseudo-3D points. These approaches replace the previous splat operation, which generates the BEV feature, with a voxel-pooling operation. This new approach voxelizes the pseudo-3D point cloud and proposes several refinement modules to enhance the 3D feature volume. OCCFormer [40] introduces a dual-path transformer block to refine the BEV slice of the 3D feature volume, enhancing the long-range modelling capability of their model. FB-OCC [41] proposes an additional backward view-transformation module to improve the semantic information in the final 3D feature volume. Multi-Scale Occ [46] leverages a multi-scale fusion mechanism to capture global and local detail information in the 3D feature volume.
While depth-estimation-based approaches have achieved remarkable performance, they have one significant drawback: requiring depth ground truth labels to boost depth-estimation performance, which in turn boosts the model’s overall performance. This requirement introduces extra effort during the training process. In this paper, we propose a method that eliminates the need for depth estimation by solely relying on multi-view images to construct a 3D feature volume. Our approach achieves superior performance compared to depth-estimation based approaches.
II-C Query-Based BEV Perception
The BEVFormer series [26, 47] is one of the pioneering query-based approaches to BEV perception. By defining a BEV query plane and utilizing intrinsic and extrinsic parameters, the query vector in each grid of the BEV plane can be projected onto multi-view images, enabling the aggregation of features and the generation of the final BEV feature. The temporal self-attention module also leverages historical BEV features to enhance the model’s performance further. Similarly, the DETR3D [48] predefines a set of 3D query points as reference points. The intrinsic and extrinsic parameters are then used to project those points onto multi-view images to aggregate features. These works successfully eliminate the need for depth estimation by employing transformer blocks and relying solely on multi-view image features for BEV perception.
II-D Query-Based 3D Semantic Occupancy Prediction
@@ III InverseMatrixVT3D
Figure 2 exhibits the overall architecture of our method. Initially, given a set of surrounding multi-camera images, we first use a 2D backbone network(e.g.ResNet101-DCN) to extract $Nc$ cameras and $L$ levels multi-scale features $X=\left\{\left\{X_{n}^{l}\right\}_{n=1}^{Nc}\in R^{C_{l}\times H_{l}\times W_{l}}\right\}_{l=1}^{L}$. For each level, we construct two projection matrices, namely, global projection matrix $VT\_XY^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l})}$ and local projection matrix $VT\_XYZ^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l}\times Z_{l})}$. The feature maps at each level are multiplied with these projection matrices, resulting in the 3D feature volume $F_{local}^{l}\in{R}^{C_{l}\times X_{l}\times Y_{l}\times Z_{l}}$ and the Bird’s-eye view (BEV) feature $F_{global}^{l}\in{R}^{C_{l}\times X_{l}\times Y_{l}}$. Subsequently, the global-local attention fusion module merges information from these two features, producing the final 3D volume. Furthermore, the 3D volume at each level is upsampled using 3D deconvolution and integrated with the higher-level 3D volume through skip-connection. Finally, the dense occupancy ground truth from [39] is applied to supervise the 3D volume at each level with a decayed loss weight.
@@ III-B Global and Local Projection Matrix Generation
Figure 3: The predefined 3D volume feature sampling process. Each voxel grid is initially divided into $N^{3}$ subspaces along the vertical and horizontal directions, and the centre of each subspace serves as a sample point. These sample points are then projected onto all the multi-view feature maps to aggregate the corresponding features for the voxel grid.
In our approach, constructing global and local projection matrices is critical in gathering information for the local 3D feature volume and the global BEV feature. This differs from the method used in Occformer[40], which relies on depth estimation and voxel pooling, or the methods employed in TPVFormer[38] and SurroundOcc[39], which involve transformer-based query. To begin with, we establish a set of predefined 3D volume spaces denoted by $F_{3D}^{l}\in{R}^{X^{l}\times Y^{l}\times Z^{l}}$ for each level of multi-view feature maps under the ego vehicle’s coordinate system. The ego vehicle is positioned at the center of these 3D volumes. Within each 3D volume, we divide the voxel grid equally into $N^{3}$ subspaces along the horizontal and vertical directions. Each center of the subspace serves as a sample point. Thus, for each level of the 3D volume space, we have a total of $X^{l}\times Y^{l}\times Z^{l}\times N^{3}$ sample points. Next, we project each sample point from each voxel grid onto the corresponding level of multi-view feature map $X^{l}$ using extrinsic and intrinsic parameters. Sample points that fall outside the boundaries of the feature maps or generate negative depth values are filtered out. The features hit by each sample point are then aggregated, resulting in a feature vector representing the corresponding voxel grid (see Figure 3).
The feature sampling process is a static mapping and can be represented by constructing the projection matrices $VT\_XYZ^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l}\times Z_{l})}$ as exhibited in Figure 4. Moreover, height information can be compressed further by constructing the projection matrix $VT\_XY^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l})}$ as shown in Figure 5. As a result, the generation of both the local 3D feature volume and global BEV feature can be greatly simplified as a matrix multiplication between the multi-view feature maps and the two projection matrices:
@@
Figure 4: The local projection matrix VT_XYZ represents the static mapping relationship for the feature sampling process of the local 3D feature volume.
Figure 5: The global projection matrix VT_XY represents the static mapping relationship for the feature sampling process of the global Bird’s Eye View (BEV) feature.
@@ III-C Global Local Attention Fusion
To enhance the ability of the final 3D feature volume to capture both global and local details, we introduce the Global-Local Attention Fusion module. The detailed structure of the Global-Local Attention Fusion module is depicted in Figure 6.
@@
Inspired by [51], the module begins by applying traditional 2D and 3D convolution operations to enhance the global BEV feature and the local 3D feature volume. Building on recent advancements that emphasize the significance of locality and efficiency in transformers [52, 53] and the importance of the BEV plane [40], we incorporate an efficient window attention module from [52] and a bottleneck ASPP module to further refine the global BEV feature, enabling it to capture long-range global semantic information. Simultaneously, the local 3D feature volume is processed through the Feed-Forward Network (FFN) to generate attention weights. The BEV feature, refined by the window attention and bottleneck ASPP module, undergoes dimension expansion and is element-wise multiplied by the attention weights. This resulting feature is then added to the refined local 3D feature volume to produce the final 3D feature volume. The whole process can be described as follows:

### v2

@@ Abstract
This paper introduces InverseMatrixVT3D, an efficient method for transforming multi-view image features into 3D feature volumes for 3D semantic occupancy prediction. Existing methods for constructing 3D volumes often rely on depth estimation, device-specific operators, or transformer queries, which hinders the widespread adoption of 3D occupancy models. In contrast, our approach leverages two projection matrices to store the static mapping relationships and matrix multiplications to efficiently generate global Bird’s Eye View (BEV) features and local 3D feature volumes. Specifically, we achieve this by performing matrix multiplications between multi-view image feature maps and two sparse projection matrices. We introduce a sparse matrix handling technique for the projection matrices to optimize GPU memory usage. Moreover, a global-local attention fusion module is proposed to integrate the global BEV features with the local 3D feature volumes to obtain the final 3D volume. We also employ a multi-scale supervision mechanism to enhance performance further. Extensive experiments performed on the nuScenes and SemanticKITTI datasets reveal that our approach not only stands out for its simplicity and effectiveness but also achieves the top performance in detecting vulnerable road users (VRU), crucial for autonomous driving and road safety. The code has been made available at: https://github.com/DanielMing123/InverseMatrixVT3D
@@ I INTRODUCTION
Understanding the surrounding scene’s three-dimensional (3D) geometry is fundamental to developing autonomous driving (AV) systems. While lidar-based methods that utilize explicit depth measurements have been performing exceptionally well on public datasets [1, 2], they are hindered by the expensive cost of sensors and the sparsity of data points. As a result, the broader application of lidar-based methods is limited.
@@
Vision-centric AV systems have garnered significant attention in recent years as a promising strategy due to its cost-effectiveness, stability, and generality. By utilizing multi-camera images as inputs, this approach has demonstrated competitive performance across various 3D perception tasks, including depth estimation [3, 4], 3D object detection [5, 6, 7], online high-definition (HD) map construction [8, 9, 10], and semantic map construction [11, 12].
Figure 1: Pipeline of three approaches: Depth-Estimation-based approach (upper), Query-based approach (middle), and our approach (lower). We simplify the generation of the 3D Feature Volume by adopting a matrix multiplication method.
@@
3D object detection based on fusing surround view cameras can be crucial in 3D perception. However, it faces challenges in handling new scenarios. One of the challenges is the finite number of semantic classes in the training dataset, making it difficult to create a model for every potential scenario that might be encountered on the road. In contrast, a more practical approach to depict the vehicle’s surrounding environment is by directly reconstructing the 3D scenes. In pursuit of this objective, several methods [13, 14, 15] have been investigated to predict the 3D occupancy of a scene directly. These methods involve voxelizing the 3D space and assigning a probability to each voxel to determine its occupancy state–whether it is occupied or not. We argue that 3D occupancy serves as an adequate representation of the vehicle’s surrounding environment. This representation inherently ensures geometric consistency and can accurately describe occluded areas. Furthermore, it is more robust towards object classes that do not exist in the training dataset. Despite the promise of these methods, their inner structure is quite complex, and some methods require additional sensors to provide supervision signals. For instance, the method proposed in [15] relies on a lidar sensor to enhance performance through depth estimation supervision (Fig. 1, upper). Additionally, the methods proposed in [13, 14] extensively employ query-based modules [16] to aggregate image features for the final 3D feature volume (Fig. 1, middle).
@@
To efficiently and effectively represent a 3D scene using 3D occupancy, we propose InverseMatrixVT3D (Fig. 1, bottom). Our method focuses on constructing projection matrices and simplifying the generation of local 3D feature volumes and global Bird’s Eye View (BEV) features through matrix multiplication between multi-scale feature maps and projection matrices. Additionally, we employ a sparse matrix handling technique to optimize GPU memory usage when using these sparse projection matrices. Furthermore, we introduce a global-local fusion module to integrate global BEV features with local 3D feature volumes, resulting in the final 3D volume. We also apply a multi-scale supervision mechanism to each level to further enhance performance. Through comparisons with other state-of-the-art (SOTA) algorithms on the nuScenes and SemanticKITTI benchmarks, we demonstrate our method not only excels in its simplicity and effectiveness but also achieves the best performance in detecting vulnerable road users (VRU), i.e. pedestrians, motorcycles, and bicycles, which is a critical task for autonomous driving and road safety.
@@
  • A novel projection matrix-based approach is proposed to simplify the local 3D feature volume and global BEV feature construction.
  • A global-local fusion module is proposed to integrate global long-range information from the BEV feature with local spatial detail information from the 3D feature volume, resulting in the final 3D volume.
  • We compare our approach with other state-of-the-art (SOTA) algorithms in the 3D semantic occupancy prediction task to prove the simplicity and effectiveness of our method.
@@ II Related Work
II-A Depth-Estimation Based 3D Semantic Occupancy Prediction
Based on the success of depth-estimation-based BEV perception algorithms, several works [17, 18, 19, 15] have focused on constructing the 3D feature volume using pseudo-3D points. These approaches replace the previous splat operation in LSS [20], which generates the BEV feature, with a voxel-pooling operation. This new approach voxelizes the pseudo-3D point cloud and proposes several refinement modules to enhance the 3D feature volume. OCCFormer [15] introduces a dual-path transformer block to refine the BEV slice of the 3D feature volume, enhancing the long-range modeling capability of their model. FB-OCC [19] proposes an additional backward view-transformation module to improve the semantic information in the final 3D feature volume. Multi-Scale Occ [18] leverages a multi-scale fusion mechanism to capture global and local detail information in the 3D feature volume.
While depth-estimation-based approaches have achieved remarkable performance, they have one significant drawback: requiring depth ground truth labels to boost depth-estimation performance, boosting the model’s overall performance. This requirement introduces extra effort during the training process. In this paper, we propose a method that eliminates the need for depth estimation by solely relying on multi-view images to construct a 3D feature volume. Our approach achieves superior performance compared to depth-estimation-based approaches.
II-B Query-Based 3D Semantic Occupancy Prediction
@@ III InverseMatrixVT3D
Fig. 2 exhibits the overall architecture of our method. Initially, given a set of surrounding multi-camera images, we first use a 2D backbone network(e.g.ResNet101-DCN) to extract $Nc$ cameras and $L$ levels multi-scale features $X=\left\{\left\{X_{n}^{l}\right\}_{n=1}^{Nc}\in R^{C_{l}\times H_{l}\times W_{l}}\right\}_{l=1}^{L}$. For each level, we construct two projection matrices, namely, global projection matrix $VT\_XY^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l})}$ and local projection matrix $VT\_XYZ^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l}\times Z_{l})}$. The feature maps at each level are multiplied with these projection matrices, resulting in the 3D feature volume $F_{local}^{l}\in\mathbb{R}^{C_{l}\times X_{l}\times Y_{l}\times Z_{l}}$ and the Bird’s-eye view (BEV) feature $F_{global}^{l}\in\mathbb{R}^{C_{l}\times X_{l}\times Y_{l}}$. Subsequently, the global-local attention fusion module merges information from these two features, producing the final 3D volume. Furthermore, the 3D volume at each level is upsampled using 3D deconvolution and integrated with the higher-level 3D volume through skip-connection. Finally, the dense occupancy ground truth from [14] is applied to supervise the 3D volume at each level with a decayed loss weight.
@@ III-B Global and Local Projection Matrix Generation
In our approach, constructing global and local projection matrices is critical in gathering information for the local 3D feature volume and the global BEV feature. This differs from the method used in Occformer[15], which relies on depth estimation and voxel pooling, or the methods employed in TPVFormer[13] and SurroundOcc[14], which involve transformer-based query.
Figure 3: The predefined 3D volume feature sampling process. Each voxel grid is initially divided into $N^{3}$ subspaces along the vertical and horizontal directions, and the center of each subspace serves as a sample point. These sample points are then projected onto all the multi-view feature maps to aggregate the corresponding features for the voxel grid.
To begin with, we establish a set of predefined 3D volume spaces denoted by $F_{3D}^{l}\in\mathbb{R}^{X^{l}\times Y^{l}\times Z^{l}}$ for each level of multi-view feature maps under the ego vehicle’s coordinate system. The ego vehicle is positioned at the center of these 3D volumes. Within each 3D volume, we divide the voxel grid equally into $N^{3}$ subspaces along the horizontal and vertical directions. Each center of the subspace serves as a sample point. Thus, for each level of the 3D volume space, we have a total of $X^{l}\times Y^{l}\times Z^{l}\times N^{3}$ sample points. Next, we project each sample point from each voxel grid onto the corresponding level of multi-view feature map $X^{l}$ using extrinsic and intrinsic parameters. Sample points that fall outside the boundaries of the feature maps or generate negative depth values are filtered out. The features hit by each sample point are then aggregated, resulting in a feature vector representing the corresponding voxel grid (see Fig. 3).
The feature sampling process is a static mapping and can be represented by constructing the projection matrices $VT\_XYZ^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l}\times Z_{l})}$ as exhibited in Fig. 4(a). Moreover, height information can be compressed further by constructing the projection matrix $VT\_XY^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l})}$ as shown in Fig. 4(b). As a result, the generation of both the local 3D feature volume and global BEV feature can be greatly simplified as a matrix multiplication between the multi-view feature maps and the two projection matrices:
@@
(a) Local projection matrix
(b) Global projection matrix
Figure 4: (a) The local projection matrix VT_XYZ represents the static mapping relationship for the feature sampling process of the local 3D feature volume. (b) The global projection matrix VT_XY represents the static mapping relationship for the feature sampling process of the global Bird’s Eye View (BEV) feature.
@@ III-C Global Local Attention Fusion
To enhance the ability of the final 3D feature volume to capture both global and local details, we introduce the Global-Local Attention Fusion module. The detailed structure of the Global-Local Attention Fusion module is depicted in Fig. 5.
@@
Inspired by [23], the module begins by applying traditional 2D and 3D convolution operations to enhance the global BEV feature and the local 3D feature volume. Building on recent advancements that emphasize the significance of locality and efficiency in transformers and the importance of the BEV plane [15], we incorporate an efficient window attention module from [24] and a bottleneck ASPP module to further refine the global BEV feature. Simultaneously, the local 3D feature volume is processed through the Feed-Forward Network (FFN) to generate attention weights. The BEV feature, refined by the window attention and bottleneck ASPP module, undergoes dimension expansion and is element-wise multiplied by the attention weights. This resulting feature is then added to the refined local 3D feature volume to produce the final 3D feature volume. The whole process can be described as follows:
@@
Method Backbone Params Resolution IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene[25] ResNet101-DCN - $200\times 200\times 16$ 23.96 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas* [26] - - $200\times 200\times 16$ 28.66 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer*[7] ResNet101-DCN 59M $200\times 200$ 30.50 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer[13] ResNet101-DCN 69M $200\times 200\times 16$ 11.51 11.66 16.14 7.17 22.63 17.13 8.83 11.39 10.46 8.23 9.43 17.02 8.07 13.64 13.85 10.34 4.90 7.37 TPVFormer* ResNet101-DCN 69M $200\times 200\times 16$ 30.86 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 C-CONet*[27] ResNet101 118M $200\times 200\times 16$ 26.10 18.40 18.60 10.00 26.40 27.40 8.60 15.70 13.30 9.70 10.90 20.20 33.00 20.70 21.40 21.80 14.70 21.30 LMSCNet*[28] - - $200\times 200\times 16$ 36.60 14.90 13.10 4.50 14.70 22.10 12.60 4.20 7.20 7.10 12.20 11.50 26.30 14.30 21.10 15.20 18.50 34.20 L-CONet*[27] - $200\times 200\times 16$ 39.40 17.70 19.20 4.00 15.10 26.90 6.20 3.80 6.80 6.00 14.10 13.10 39.70 19.10 24.00 23.90 25.10 35.70 SurroundOcc*[14] ResNet101-DCN 180M $200\times 200\times 16$ 31.49 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 InverseMatrixVT3D* (Base) ResNet101-DCN 67M $200\times 200\times 16$ 31.85 18.88 18.39 12.46 26.30 29.11 11.00 15.74 14.78 11.38 13.31 21.61 36.30 19.97 21.26 20.43 11.49 18.47 InverseMatrixVT3D* (Post-Fix) ResNet101-DCN 67M $200\times 200\times 16$ 26.79 15.81 16.47 10.27 21.28 28.29 8.32 13.29 12.90 8.41 10.96 18.49 32.43 11.79 18.27 15.45 9.87 16.52 InverseMatrixVT3D* (Pre-Fix) ResNet101-DCN 67M $200\times 200\times 16$ 31.30 18.42 18.06 12.72 25.99 28.00 10.15 15.98 14.31 10.61 12.49 20.58 35.61 19.40 21.00 20.30 11.06 18.59
TABLE I: 3D semantic occupancy prediction results on nuScenes validation set. * means method is trained with dense occupancy labels from [14].
Method IoU mIoU road sidewalk parking other-ground building car truck bicycle motorcycle other-vehicle vegetation trunk terrain person bicyclist motorcyclist fence pole traf.sign LMSCNet [28] 28.61 6.70 40.68 18.22 4.38 0.00 10.31 18.33 0.00 0.00 0.00 0.00 13.66 0.02 20.54 0.00 0.00 0.00 1.21 0.00 0.00 AICNet [29] 29.59 8.31 43.55 20.55 11.97 0.07 12.94 14.71 4.53 0.00 0.00 0.00 15.37 2.90 28.71 0.00 0.00 0.00 2.52 0.06 0.00 3DSketch [30] 33.30 7.50 41.32 21.63 0.00 0.00 14.81 18.59 0.00 0.00 0.00 0.00 19.09 0.00 26.40 0.00 0.00 0.00 0.73 0.00 0.00 JS3C-Net [31] 38.98 10.31 50.49 23.74 11.94 0.07 15.03 24.65 4.41 0.00 0.00 6.15 18.11 4.33 26.86 0.67 0.27 0.00 3.94 3.77 1.45 MonoScene [25] 36.86 11.08 56.52 26.72 14.27 0.46 14.09 23.26 6.98 0.61 0.45 1.48 17.89 2.81 29.64 1.86 1.20 0.00 5.84 4.14 2.25 TPVFormer [13] 35.61 11.36 56.50 25.87 20.60 0.85 13.88 23.81 8.08 0.36 0.05 4.35 16.92 2.26 30.38 0.51 0.89 0.00 5.94 3.14 1.52 InverseMatrixVT3D (Base) 36.22 11.81 52.99 25.84 20.04 0.09 13.17 24.08 10.25 1.85 2.65 6.80 16.98 3.09 27.77 4.01 3.13 0.00 4.94 4.05 2.67
TABLE II: 3D semantic scene completion performance on SemanticKITTI validation set.